OpenStudy (anonymous):

I need help answering some word problems which freezes first ? how much colder does it have to be for both to freeze? ... oxygen is -218.79 c and hydrogen is -252.8 c

OpenStudy (sriram):

i answered the same question earlier some time back!!

OpenStudy (anonymous):

the thing was glitching on me

OpenStudy (anonymous):

open study was having problems it was glitching freezing srry

OpenStudy (sriram):

look -218>-252 hence oxygen freezes at higher temp oxygen freezes first

OpenStudy (anonymous):

so how much colder does it have to be for both to freeze

OpenStudy (sriram):

when temp is -218 only oxygen freezes and when u keep lowerin the temp to -252even hydrogen freezes both freeze at -252

OpenStudy (anonymous):

ohh ok could u help me with easier ones ?

OpenStudy (sriram):

post them

OpenStudy (anonymous):

kk hold up

OpenStudy (anonymous):

1. How many degrees did the temperature drop? ... at 3pm is 1 c the temp drop to -9c at 3 am in quebec

OpenStudy (sriram):

temp drop =initial temp-final temp = 1- (-9) =10c convert it to quebec

OpenStudy (anonymous):

2. How much is the drop in temperature from day to night ? .... on the side of the moon that the sun is shining on the temp can reach 260 f and on the dark side it gets as cold as 280 f. 3. how much cooler is the melting point of silver than gold? gold melts at 1946 f and silver melts at 1762.

OpenStudy (sriram):

the 2nd i think u switched the values coz in your question, the day is colder than the night!

OpenStudy (anonymous):

do u want me to readu the whole ?

OpenStudy (sriram):

no just check whether day temp< night temp

OpenStudy (anonymous):

day temp is 260 far and night is -280 f

OpenStudy (sriram):

k fine then temp drops by 260+280=540f this is bcoz initially it was at 260 then to reach -280 it has to drop by 540 f

OpenStudy (anonymous):

ok

OpenStudy (sriram):

the 3rd just subtract 1946-1762=184f

OpenStudy (anonymous):

I have more ? but idk if u want to do more its up to you.

OpenStudy (anonymous):

these are short ?

OpenStudy (sriram):

well will try tomm at the same time my moms buggin me for dinner goodnight pal.

OpenStudy (anonymous):

thanks

OpenStudy (sriram):

yep these are short